OpenStudy (anonymous):

Find the area of the triangle formed by 3 points on the curve y=x^4 -2x^2 +4. where the slope of the tangent is 0. I am completely lost as to how to solve this question. Any help is appreciated :)

OpenStudy (anonymous):

if you use a graphing calculator you can see it looks like a W, where the end points are (-1, 3), (0, 4), and (1, 3). Therefore, the base is 2 and the height is 1. 1/2bh = 1/2(2)(1) = 1

OpenStudy (anonymous):

whoops, didn't know you weren't supposed to give direct answers. Hopefully I made a mistake...check it so you can run through the process and be sure!